OpenStudy (anonymous):

6(4/5)n-(8/9)=(7/15)n

OpenStudy (mertsj):

\[6(\frac{4}{5})n-\frac{8}{9}=\frac{7}{15}n\]

OpenStudy (mertsj):

\[\frac{24}{5}n-\frac{8}{9}=\frac{7}{15}n\] \[-\frac{8}{9}=\frac{7}{15}n-\frac{24}{5}n\] \[\frac{-8}{9}=\frac{-13}{3}n\] \[n=\frac{8}{39}\]